OpenStudy (anonymous):

which of the following are identities? a. sin8x = 2sin4x cos4x b.(sinx-cosx) ^2 = 1 + sin2x c. (1-tan^2x)/2tanx = 1/ tan2x d. (sinx +sin5x)/ (cosx+ cos5x) = tan3x

OpenStudy (baldymcgee6):

just put in a value for x and see if it is true.

OpenStudy (anonymous):

i can't do c

OpenStudy (anonymous):

whenever i try to input values it never works

OpenStudy (anonymous):

.

OpenStudy (anonymous):

that was helpful

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

can you please help me ?

OpenStudy (anonymous):

The answer can't B (sinx - cosx)^2 = (sinx)^2 - 2sinxcosx + (cosx)^2 (sinx)^2 + (cosx)^2 = 1; 2sinxcosx = sin2x therefore, (sinx - cosx)^2 = 1- sin2x

OpenStudy (anonymous):

Answer is A http://www.wolframalpha.com/input/?i=+sin8x+%3D+2sin4x+cos4x

OpenStudy (anonymous):

a ..... not b ..okay what about c and D?

OpenStudy (raden):

c. correct to

OpenStudy (raden):

d. correct